### v1

@@ Abstract
Carbon-based 2D heterostructures represent an attractive platform for nanoelectronics owing to their tunable electronic and transport properties, yet achieving precise control over their fabrication remains elusive. Here, we demonstrate the on-surface synthesis of covalently bonded lateral heterostructures between armchair graphene nanoribbons and metalated hydrogenated graphdiyne networks on Au(111). Atomic-resolution scanning tunnelling microscopy combined with density functional theory reveals the formation mechanism of covalent interfacial bonds and highlights the critical influence of surface chemistry. In particular, chemisorbed bromine atoms suppress junction formation, while controlled atomic hydrogen dosing increases the bonding efficiency to 71%. Electronic structure and transport calculations demonstrate how the metallic substrate influences the supported heterostructure, whereas in the freestanding limit, the two carbon subsystems retain their intrinsic properties, forming an atomically narrow junction that enables voltage-tunable spatial current separation. These results define a viable strategy for engineering graphene–graphdiyne heterostructures and advance the design of all-carbon nanoscale electronic architectures.
@@
2D carbon heterostructures, On-surface synthesis, Graphdiyne-Graphene interfaces, Carbon nanoelectronics, Density functional theory, Transmission functionMain
@@
Low coverage aGNRs samples (0.2-0.3 ML) are realized as scaffolds to host the growth of metalated hGDY in-between the nanoribbons: once the GNRs are obtained [40, 24, 41], tBEB molecules are deposited on the surface decorated with the GNRs (see Figures S1-S2 and Methods). Upon annealing at 400 K to promote the ordering of the metalated hGDY, Br atoms coming from the dehalogenation of the hGDY precursor are found mainly along the edges of the aGNRs, being stabilized by the interaction with the H atoms saturating the nanoribbons (see Figure 1c-e and Methods).
According to our interpretation, following this annealing step, Au adatoms terminating the metalated hGDY tend to form complexes with Br atoms decorating the nanoribbons (see Figure 1c,e), as confirmed by ab initio calculations (see S3), preventing the formation of the hGDY–aGNR heterojunction.
@@
Based on the total energy differences reported in Table 1, we can infer that the formation process of the hGDY–aGNR heterojunction may take place through the initial formation of a “Top” configuration, i.e. the most likely due to the proximity of the ending carbon bond in the hGDY fragment and the aGNR, possibly followed by a distortion of the acetylene bond that drives the system in configuration “1H” or “2H”.
Indeed, the formation energy for the latter, calculated as detailed in Methods, is equal to -3.55 eV.
@@ Acknowledgments
A.C., A.L.B. and C.S.C acknowledge funding by: - Funder: project funded under the National Recovery and Resilience Plan (NRRP), Mission 4 Component 2 Investment 1.3 - Call for tender No. 341 of 15.03.2022 of Ministero dell’Universitá e della Ricerca (MUR); funded by the European Union NextGenerationEU - Award Number: project code PE0000021, Concession Decree No. 1561 of 11.10.2022 adopted by Ministero dell’Universitá e della Ricerca (MUR), CUP D43C22003090001, Project title “Network 4 Energy Sustainable Transition NEST’ ’. F.S. acknowledges the financial support from the University of Padova through the grant P-DiSC#02BIRD2024-UNIPD VS-PpyTM. J.L.C. acknowledges financial support from the Spanish Ministry of Science and Innovation (Grant PID2022_138750NB-C21 and “Severo Ochoa’ ’ Programme for Centres of Excellence in R&D CEX2023-001286-S funded by MCIN/AEI/ 10.13039/501100011033 and by ERDF “A way of making Europe’ ’ ), from the regional Governments of Aragon (E12_23R) and from the European Commission (project ULTIMATE-I, grant ID 101007825). J.L.C. and A.E.C. further acknowledge the use of Servicio General de Apoyo a la Investigación-SAI and the Laboratorio de Microscopías Avanzadas of the Universidad de Zaragoza. S.A. acknowledges the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources and support, and the financial support from INFN (Progetto iniziativa specifica “TIME2QUEST”). S.A. also acknowledges Federico Civaia for performing preliminary calculations on systems similar to those presented in the manuscript. A.O.B. acknowledges support from the Ministero dell’Universitá e della Ricerca (MUR) and the University of Pavia through the program “Dipartimenti di Eccellenza” 2023–2027.

### v2

@@ Abstract
Two-dimensional heterostructures combining sp-sp2 hybridization—blending graphene with graphyne-based allotropes—offer substantial potential for enhancing the tunability of electronic and transport properties while providing significant structural flexibility. These attributes are desirable for next generation nanoscale electronic applications. Despite such potential, their experimental realization remains elusive, as synthesized carbon heterostructures are limited to doped, graphene-based systems exhibiting exclusively sp2 hybridization. Here, we demonstrate the on-surface synthesis of covalently bonded sp-sp2 lateral heterostructures between graphene nanoribbons and graphdiyne networks on Au(111). Atomic-resolution scanning tunnelling microscopy, combined with density functional theory, reveals the formation mechanism of the covalent interfacial bonds between nanoribbons and graphdiynes, also highlighting the key role of surface chemistry. Bromine atoms deriving from the molecules dehalogenation and chemisorbed along the nanoribbon inhibit the junction formation, but bonding efficiency can be boosted up to 71% by controlled removal of these by-products. Electronic structure and transport calculations show that the 2D heterostructure by itself is characterized by disentangled properties for the two subsystems, forming an atomically narrow junction enabling voltage-tunable spatial current separation in two dimensions. There results define a viable strategy for engineering graphene-based sp-sp2 heterostructures, paving the way for the design and synthesis of all-carbon nanoscale electronic architectures.
@@
2D sp-sp2 carbon heterostructures, On-surface synthesis, Graphdiyne-Graphene interfaces, Carbon nanoelectronics, Density functional theory, Transmission functionIntroduction
@@
Low coverage aGNRs samples (0.2-0.3 ML) are realized as scaffolds to host the growth of metalated hGDY in-between the nanoribbons: once the GNRs are obtained [40, 24, 41], tBEB molecules are deposited on the surface decorated with the GNRs (see Figures S1-S2 and Methods). Upon annealing at 400 K to promote the ordering of the metalated hGDY, Br atoms coming from the dehalogenation of the hGDY precursor are found mainly along the edges of the aGNRs, being stabilized by the interaction with the H atoms saturating the nanoribbons (see Figure 1c-e and Methods). According to our interpretation, following this annealing step, Au adatoms terminating the metalated hGDY tend to form complexes with Br atoms decorating the nanoribbons (see Figure 1c,e), as confirmed by ab initio calculations (see S3), preventing the formation of the hGDY–aGNR heterojunction.
@@
Based on the total energy differences reported in Table 1, we can infer that the formation process of the hGDY–aGNR heterojunction may take place through the initial formation of a “Top” configuration, i.e. the most likely due to the proximity of the ending carbon bond in the hGDY fragment and the aGNR, possibly followed by a distortion of the acetylene bond that drives the system in configuration “1H” or “2H”. Indeed, the formation energy for the latter, calculated as detailed in Methods, is equal to -3.55 eV.
@@ Acknowledgments
A.C., A.L.B. and C.S.C acknowledge funding by: Funder: project funded under the National Recovery and Resilience Plan (NRRP), Mission 4 Component 2 Investment 1.3 - Call for tender No. 341 of 15.03.2022 of Ministero dell’Universitá e della Ricerca (MUR); funded by the European Union NextGenerationEU - Award Number: project code PE0000021, Concession Decree No. 1561 of 11.10.2022 adopted by Ministero dell’Universitá e della Ricerca (MUR), CUP D43C22003090001, Project title “Network 4 Energy Sustainable Transition NEST’ ’. F.S. acknowledges the financial support from the University of Padova through the grant P-DiSC#02BIRD2024-UNIPD VS-PpyTM. J.L.C. acknowledges financial support from the Spanish Ministry of Science and Innovation (Grant PID2022_138750NB-C21 and “Severo Ochoa’ ’ Programme for Centres of Excellence in R&D CEX2023-001286-S funded by MCIN/AEI/ 10.13039/501100011033 and by ERDF “A way of making Europe’ ’ ), from the regional Governments of Aragon (E12_23R) and from the European Commission (project ULTIMATE-I, grant ID 101007825). J.L.C. and A.E.C. further acknowledge the use of Servicio General de Apoyo a la Investigación-SAI and the Laboratorio de Microscopías Avanzadas of the Universidad de Zaragoza. S.A. acknowledges the CINECA award under the ISCRA initiative , for the availability of high-performance computing resources and support, and the financial support from INFN (Progetto iniziativa specifica “TIME2QUEST”). S.A. also acknowledges Federico Civaia for performing preliminary calculations on systems similar to those presented in the manuscript. A.O.B. acknowledges support from the Ministero dell’Universitá e della Ricerca (MUR) and the University of Pavia through the program “Dipartimenti di Eccellenza” 2023–2027.